OpenStudy (anonymous):

solve the following equation in(2x + 1) = 10 ??? how do i do this

OpenStudy (anonymous):

Is that a natural logarithm?

OpenStudy (anonymous):

idk

hero (hero):

I think he means ln(2x + 1) = 10

OpenStudy (anonymous):

yea

hero (hero):

The L looks Iike an i

OpenStudy (anonymous):

oh sorry

hero (hero):

That's the first time I've seen anyone do that.

OpenStudy (anonymous):

do you know how to solve it?

hero (hero):

Yup

OpenStudy (anonymous):

(e^10 - 1)/2

hero (hero):

@MrMoose, posting that alone doesn't provide much of an explanation for @tutorme

OpenStudy (anonymous):

sorry I joined a few minutes ago

OpenStudy (anonymous):

any way

OpenStudy (anonymous):

can you explain how you got that cause i don't know how you started

OpenStudy (anonymous):

raise both sides to the e power to get rid of ln subtract both sides by 1 divide both sides by 2

OpenStudy (anonymous):

a logarithm is just a measure of what you have to raise something to to get that number

OpenStudy (anonymous):

the number inside the parenthesis

OpenStudy (anonymous):

so raising e to the natural log of something by definition gives you that something

hero (hero):

I see what the problem is now @MrMoose. Ideally, it would be best if you began with some general formula when explaining these.

OpenStudy (anonymous):

so e^ln(u) = u where u is some expression

OpenStudy (anonymous):

@Hero what is the formula?

hero (hero):

In this case, the general formula would be \(\ln_{e}(y) = x\) is equivalent to \(e^x= y\)

hero (hero):

By default if the index isn't given, then it is assumed to be e

hero (hero):

In this case, y = (2x + 1), x = 10

hero (hero):

Inserting those values into the formula we get: \(\ln_e(2x+1) = 10\) which becomes \(e^{10} = 2x + 1\) after substitution.

hero (hero):

From there, you solve for x to get: \(\frac{e^{10} - 1}{2} = x\)

OpenStudy (anonymous):

ok i get it a little.. can you help me with another one

OpenStudy (anonymous):

e^2x - 2e^x -3=0

OpenStudy (anonymous):

just one last one

OpenStudy (anonymous):

@Hero